# PreGIP: Watermarking the Pretraining of Graph Neural Networks for Deep Intellectual Property Protection

### v1

@@ Abstract
Pretraining on Graph Neural Networks (GNNs) has shown great power in facilitating various downstream tasks. As pretraining generally requires huge amount of data and computational resources, the pretrained GNNs are high-value Intellectual Properties (IP) of the legitimate owner. However, adversaries may illegally copy and deploy the pretrained GNN models for their downstream tasks. Though initial efforts have been made to watermark GNN classifiers for IP protection, these methods require the target classification task for watermarking, and thus are not applicable to self-supervised pretraining of GNN models. Hence, in this work, we propose a novel framework named PreGIP to watermark the pretraining of GNN encoder for IP protection while maintain the high-quality of the embedding space. PreGIP incorporates a task-free watermarking loss to watermark the embedding space of pretrained GNN encoder. A finetuning-resistant watermark injection is further deployed. Theoretical analysis and extensive experiments show the effectiveness of PreGIP in IP protection and maintaining high-performance for downstream tasks.
Machine Learning, ICML
1 Introduction
@@
Graph Neural Networks (GNNs) have shown great power in modeling graphs for various tasks. However, the success of GNNs rely on sufficient high-quality labels, which can be difficult to collect for the target downstream tasks. To address this problem, various self-supervised learning methods such as Graph Contrastive Learning (GCL) (You et al., 2020) have been proposed to pretrain GNN models to facilitate the training on downstream tasks. These pretraining methods have been widely applied for various domains such as molecular property prediction (Hu et al., 2020) and protein analysis (Zhang et al., 2022).
@@
The pretraining of GNN models generally demands massive computation on large-scale dataset. For example, self-supervised pretraining in chemistry (Hu et al., 2020) takes over 24 hours on an A100 GPU. Similarly, pretraining on proteins (Zhang et al., 2022) is conducted on datasets with around 1M protein structures predicted by AlphaFold2. Consequently, pretrained GNN models are high-value intellectual properties (IP) required to be protected. In practice, the pretrained models are often shared with authorized users for the downstream classifier training. However, this poses a significant risk that unauthorized individuals may steal and deploy the pretrained GNN models for their own tasks without the model owner’s permission. Therefore, it is crucial to protect the IP of pretrained GNN models.
@@
Several model ownership verification approaches (Adi et al., 2018; Yang et al., 2021; Xu et al., 2023a; Zhao et al., 2021) have been proposed to protect the IP of GNN classifiers trained in a supervised manner. A typical way of IP protection is to add watermark during the training of GNN classifiers. As illustrated in Fig. 1, during the training phase, watermarking approaches will force the protected model to predict trigger samples as certain pre-defined classes (Xu et al., 2023a; Adi et al., 2018). As a result, the predictions on the trigger samples can work as IP message for model ownership verification. Specifically, a suspect model that gives identical predictions as the protected model on the trigger samples can be identified as piracy model stolen from the protected model. However, the existing watermarking methods are not applicable for pretraining GNNs. This is because pretraining of GNNs only use pretext tasks to obtain a high-quality embedding space. As downstream tasks and labels are generally unknown at the stage of pretraining, model owner cannot utilize trigger sample classification in downstream tasks to convey IP message.
@@
Therefore, in this paper, we propose a task-free IP protection framework which aims to watermark the representation space for pretraining GNN encoder. Specifically, as shown in Fig. 2, we implement watermarking during pretraining by ensuring similar representations for each pair of dissimilar watermark graphs. If the representations of two paired watermark graphs are close enough, they will receive the same predictions from the downstream classifier under mild conditions, which is verified by our theoretical analysis in Theorem 3.1. By contrast, an independent model trained on non-watermarked data would obtain distinct/dissimilar representations for each watermarked graph pair, resulting in inconsistent predictions. Hence, as shown in Fig. 2, during the verification, we can identify the piracy model by assessing the consistency of its predictions on downstream tasks for each pre-defined watermark graph pair.
Figure 1: GNN classifier watermarking and verification.
@@
However, watermarking the embedding space during unsupervised pretraining of GNNs to achieve the above goal is a non-trivial task. There are two major challenges to be resolved. Firstly, as mentioned above, we aim to ensure similar representations for paired watermark graphs, which will obtain same predictions from downstream classifiers as the IP message. How can we achieve this and simultaneously preserve the high quality of the embedding space for non-watermark graphs? Secondly, to improve the performance on downstream tasks, unauthorized adversaries often finetune the pretrained GNN encoder. The fine-tuning process may erase the watermarks injected to the pretrained GNN model. In an attempt to address the challenges, we propose a novel Pretrained GNN IP protection framework (PreGIP). Specifically, PreGIP proposes a task-free watermarking loss to guarantee both watermarking performance and discriminability of graph representations. In addition, to reduce the effects of watermarking to real graphs, watermarking graph pairs are consist of synthetic graphs. And a finetuning-resistant watermarking approach is deployed in PreGIP to ensure the identification of suspect model built from finetuning the protected GNN encoder on various tasks. The effectiveness of PreGIP is demonstrated theoretically and empirically. In summary, our main contributions are:
@@
  • We study a new problem of watermarking the pretraining of GNN encoder without the downstream task information for deep IP protection.
@@
  • We propose a novel framework PreGIP that can effectively watermark the pretraining of GNN and resists to the finetuning of the pretrained GNN.
@@
Figure 2: Overview framework of our PreGIP.
2 Problem Formulation
Preliminaries of Model Watermarking. Model watermarking is a IP protection technique to identify the piracy model. Specifically, a piracy model is an unauthorized copy or finetuned version of a protected AI model, exploited by adversaries for their specific purposes. The model watermarking consists of two stages, i.e., watermark embedding and watermark verification. In the watermark embedding phase, unique secret keys $\mathcal{K}$ and corresponding IP message $\mathcal{I}$ will be built and injected to the deep model $f$. As Fig. 1 shows, a typical form of secret keys and IP message for supervised GNN classifier are special input samples and corresponding model predictions. When verifying the ownership of model ${f}$, model owner will extract the IP message with $\mathcal{K}$ by ${\mathcal{I}}=ExtractIP({f},\mathcal{K})$. A piracy model will retrain an IP message similar to the protected model, while an independent model will not.
Threat Model. In this paper, we focus on the IP protection of pretrained GNN encoders. Adversaries will legitimately deploy the protected GNN encoder $f_{E}$ to obtain the classifier ${f}_{A}:\mathcal{G}\rightarrow y$ for their target classification tasks. Parameters of protected GNN encoders may be finetuned by adversaries. The target tasks of adversaries are unknown for the model owner during the watermark embedding phase. During the verification phase, we assume a black-box setting, where the model owner can only obtain the predictions of suspect model on queried samples. This is reasonable because adversaries are typically reluctant to disclose models constructed without authorization.
Problem Definition. In this paper, we focus on pretraining for graph-level tasks. Let $\mathcal{D}=\{\mathcal{G}_{i}\}_{i=1}^{|\mathcal{D}|}$ denote the set of unlabeled graphs used for GNN encoder pretraining. Given the above description, the problem of watermarking the pretraining of GNNs can be formulated by:
@@ 2 Problem Formulation
Given a training set $\mathcal{D}$, we aim to construct and embed $(\mathcal{K},\mathcal{I})$, i.e., secret keys and IP message, into the pretraining of GNN encoder $f_{E}:\mathcal{G}\rightarrow\mathbb{R}^{d}$. The watermarked GNN encoder $f_{E}$ should produce high-quality embeddings effective in downstream tasks. And when adversaries deploy and finetune watermarked GNN encoder $f_{E}$ to build unauthorized classifier ${f}_{A}:\mathcal{G}\rightarrow y$, the IP message can be preserved and extracted from the piracy model ${f}_{A}$ by $\mathcal{I}=ExtractIP({f}_{A},\mathcal{K})$. ${f}_{A}$ is composed of a classifier $f_{C}$ on top of $f_{E}$.
3 Methodology
@@
In this section, we give the details of our proposed PreGIP, which is illustrated in Fig. 2. As shown in Fig. 2, PreGIP aims to enforce similar representations for multiple pairs of watermark graphs. Then, downstream classifiers topped on the pretrained GNN encoder will give same predictions on each pre-defined watermark graph pair. As a result, the consistency of predictions on the paired watermark graphs can indicate whether the suspect classifier is constructed from the watermarked GNN encoder. There are two technical challenges remain to be addressed: (i) how can we watermark the representation space of the pretrained GNN encoder and simultaneously preserve the high quality of the learned representation space? (ii) adversaries may finetune the pre-trained GNN for various downstream tasks. How to ensure watermark is preserved after finetuning the pretrained GNN encoder by adversaries? To address the above challenges, PreGIP incorporates a task-free watermarking loss that guarantees both watermarking performance and discriminability of graph embeddings. A finetuning-resistant watermark injection algorithm is also deployed in PreGIP. Next, we give the details of each component.
3.1 Task-Free Watermarking Framework
@@
Overview of Framework. Since adversaries’ downstream tasks are generally unavailable for the pretrained GNN owner, we propose a task-free watermarking framework. As presented in Fig. 2, multiple pairs of watermark graphs $\{(\mathcal{G}_{w}^{a},\mathcal{G}_{w}^{b})\}_{w=1}^{|\mathcal{K}|}$ are constructed as secret keys $\mathcal{K}$ for IP protection. During the watermark embedding phase, the proposed task-free watermark loss will ensure highly similar representations for each pair of watermark graphs $(\mathcal{G}_{w}^{a},\mathcal{G}_{w}^{b})$, while preserving the quality of representation space. In Theorem 3.1, we theoretically show that similar representations will yield same predictions from the downstream classifier. Thus, the consistency of predictions from the suspect GNN classifier $f_{A}$ on each pair of $(\mathcal{G}_{w}^{a},\mathcal{G}_{w}^{b})$ can work as IP message to identify unauthorized deployment of watermarked GNN encoder. Next, we present the detailed implementation of the task-free watermarking.
@@
Watermark Graph Pair Construction. To watermark pretraining of GNN encoder, PreGIP will enforce similar representations on distinct graphs, leading to identical predictions in case of piracy model. One straightforward way of construction is to randomly sample real graphs as watermark graph pairs. However, enforcing close embeddings for distinct real graphs may significantly degrade the embedding quality on real-world test samples. Therefore, we utilize synthetic graphs to construct the watermark graph pair set $\mathcal{K}=\{(\mathcal{G}_{w}^{a},\mathcal{G}_{w}^{b})\}_{w=1}^{|\mathcal{K}|}$. Specifically, the node attributes $\mathbf{x}$ of each graph in $\mathcal{K}$ are sampled by:
$\mathbf{x}\sim\mathcal{N}(\bf{\mu},\bf{\sigma}^{2}),$ (1)
@@
where $\mu$ and $\sigma$ denote the mean and standard deviation of the node attributes in the pretraining dataset $\mathcal{D}$, respectively. As for the graph structure generation, we use a Erdos-Renyi (ER) random graph model (ERDdS & R&wi, 1959) which independently connects each pair of nodes with probability $p$. For each pair of watermark graphs $(\mathcal{G}_{w}^{a},\mathcal{G}_{w}^{b})$, the probability parameter $p$ and graph size will be set as different values to make $\mathcal{G}_{w}^{a}$ and $\mathcal{G}_{w}^{b}$ distinct. In this way, $\mathcal{G}_{w}^{a}$ and $\mathcal{G}_{w}^{b}$ will receive dissimilar predictions from independently trained models trained on non-watermarked datasets.
@@
Watermarking Loss. Apart from the term for high similarity between paired watermark graphs, we also incorporate a loss term to encourage that embeddings of synthetic watermark graphs have no overlap with those of real-world graphs. This will offer two advantages over the loss that solely enforces similar embeddings between watermark graph pairs: (i) it prevents the trivial solution that $f_{E}$ outputs the same embeddings to all graphs; (ii) embedding space of synthetic watermark graphs will remain distinct from that of real-world graphs, thereby preserving the quality of real graph embeddings for downstream tasks.
@@
Secifically, let $\mathbf{h}_{w}^{a}$ and $\mathbf{h}_{w}^{b}$ denote the graph representations of each pair of watermark graphs $(\mathcal{G}_{w}^{a},\mathcal{G}_{w}^{b})$ generated by $f_{E}$, the watermarking loss is formulated as:
$\min_{\theta}$ $\quad\mathcal{L}_{W}(\theta)=\sum_{(\mathcal{G}_{w}^{a},\mathcal{G}_{w}^{b})\in\mathcal{K}}\Big{(}\|\mathbf{h}_{w}^{a}-\mathbf{h}_{w}^{b}\|^{2}+$ (2)
@@
$\min_{\theta}\quad\mathcal{L}_{Pre}(\theta)+\lambda\mathcal{L}_{W}(\theta),$ (3)
@@
Ownership Verification. For a piracy classifier that utilizes the watermarked GNN encoder $f_{E}$, it will be more likely to give identical predictions on each pair of $(\mathcal{G}_{w}^{a},\mathcal{G}_{w}^{b})$; while an independently trained classifier would not exhibit this behavior. Thus, we extract the IP message by computing the portion of same predictions on watermark graph pair set $\{(\mathcal{G}_{w}^{a},\mathcal{G}_{w}^{b})\}_{w=1}^{|\mathcal{K}|}$. Given a GNN classifier $f_{A}:\mathcal{G}\rightarrow y$ to be verified, its IP indication score for ownership verification is:
$I(f_{A})=\frac{\sum_{w=1}^{|\mathcal{K}|}\mathbbm{1}(f_{A}(\mathcal{G}_{w}^{a})=f_{A}(\mathcal{G}_{w}^{b}))}{|\mathcal{K}|},$ (4)

### v2

# PreGIP: Watermarking the Pretraining of Graph Neural Networks for Deep IP Protection

@@ Abstract.
Pretraining on Graph Neural Networks (GNNs) has shown great power in facilitating various downstream tasks. As pretraining generally requires huge amount of data and computational resources, the pretrained GNNs are high-value Intellectual Properties (IP) of the legitimate owner. However, adversaries may illegally copy and deploy the pretrained GNN models for their downstream tasks. Though initial efforts have been made to watermark GNN classifiers for IP protection, these methods are not applicable to self-supervised pretraining of GNN models. Hence, in this work, we propose a novel framework named PreGIP to watermark the pretraining of GNN encoder for IP protection while maintaining the high-quality of the embedding space. PreGIP incorporates a task-free watermarking loss to watermark the embedding space of pretrained GNN encoder. A finetuning-resistant watermark injection is further deployed. Theoretical analysis and extensive experiments show the effectiveness of PreGIP. The code can be find in https://anonymous.4open.science/r/PreGIP-semi/ and https://anonymous.4open.science/r/PreGIP-transfer/.
Graph Neural Networks; Pretraining; IP Protection
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737089isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Machine learning
1. Introduction
@@
Graph Neural Networks (GNNs) have shown great power in modeling graphs for various tasks (Xu et al., 2024; Dai et al., 2021b; Dai and Wang, 2021a, 2022). Recently, inspired by the success of pretraining on large-scale text data  (Ouyang et al., 2022; Lin et al., 2024, 2025b; Wu et al., 2025), various self-supervised learning methods such as Graph Contrastive Learning (GCL) (You et al., 2020) have been proposed to pretrain GNN models to facilitate the adoption of GNNs on downstream tasks. These pretraining methods have been widely applied for various domains such as molecular property prediction (Hu et al., 2020; Lin et al., 2025a) and protein analysis (Zhang et al., 2022), and energy network anomaly detection (Dai and Chen, 2022).
@@
The pretraining of GNN models generally demands massive computation on large-scale datasets. For example, self-supervised pretraining in chemistry (Hu et al., 2020) takes over 24 hours on an A100 GPU. Similarly, pretraining on proteins (Zhang et al., 2022; Li et al., 2025) is conducted on datasets with around 1M protein structures predicted by AlphaFold2. Consequently, pretrained GNN models are high-value intellectual properties (IP) required to be protected. In practice, the pretrained models are often shared with authorized users for downstream classifier training. However, this poses a significant risk that unauthorized individuals may steal and deploy the pretrained GNN models for their own tasks without the model owner’s permission. Therefore, it is crucial to protect the IP of pretrained GNN models.
@@
Several model ownership verification approaches (Adi et al., 2018; Yang et al., 2021; Xu et al., 2023; Zhao et al., 2021) have been proposed to protect the IP of GNN classifiers trained in a supervised manner. A typical way of IP protection is to add watermark during the training of GNN classifiers. As illustrated in Fig. 1, during the training phase, watermarking approaches will force the protected model to predict trigger samples as certain pre-defined classes (Xu et al., 2023; Adi et al., 2018). As a result, the predictions on the trigger samples can work as an IP message for model ownership verification. Specifically, a suspect model that gives identical predictions as the protected model on the trigger samples can be identified as a piracy model stolen from the protected model. However, the existing watermarking methods are not applicable for pretraining GNNs. This is because pretraining of GNNs only uses pretext tasks to obtain a high-quality embedding space. As downstream tasks and labels are generally unknown at the stage of pretraining, the model owner cannot utilize trigger sample classification in downstream tasks to convey IP messages.
@@
Therefore, in this paper, we propose a task-free IP protection framework that aims to watermark the representation space for pretraining GNN encoder. Specifically, as shown in Fig. 2, we implement watermarking during pretraining by ensuring similar representations for each pair of dissimilar watermark graphs. If the representations of two paired watermark graphs are close enough, they will receive the same predictions from the downstream classifier under mild conditions, which is verified by our theoretical analysis in Theorem 3.1. By contrast, an independent model trained on non-watermarked data would obtain distinct/dissimilar representations for each watermarked graph pair, resulting in inconsistent predictions. Hence, as shown in Fig. 2, during the verification, we can identify the piracy model by assessing the consistency of its predictions on downstream tasks for each pre-defined watermark graph pair.
@@
However, watermarking the embedding space during unsupervised pretraining of GNNs to achieve the above goal is a non-trivial task. There are two major challenges to be resolved. Firstly, as mentioned above, we aim to ensure similar representations for paired watermark graphs, which will obtain the same predictions from downstream classifiers as the IP message. How can we achieve this and simultaneously preserve the high quality of the embedding space for non-watermark graphs? Secondly, to improve the performance on downstream tasks, unauthorized adversaries often finetune the pretrained GNN encoder. The fine-tuning process may erase the watermarks injected into the pretrained GNN model. In an attempt to address the challenges, we propose a novel Pretrained GNN IP protection framework (PreGIP). Specifically, PreGIP proposes a task-free watermarking loss to guarantee both watermarking performance and discriminability of graph representations. In addition, to reduce the effects of watermarking on real graphs, watermarking graph pairs consist of synthetic graphs. And a finetuning-resistant watermarking approach is deployed in PreGIP to ensure the identification of suspect models built from finetuning the protected GNN encoder on various tasks. The effectiveness of the proposed PreGIP is demonstrated theoretically and empirically. In summary, our main contributions are:
@@
  • We study a new problem of watermarking the pretraining of GNN without the downstream task information for deep IP protection.
@@
  • We propose a novel framework that can watermark the pretraining of GNN and resist the finetuning of the pretrained GNN.
@@
2. Problem Formulation
Figure 1. GNN classifier watermarking and verification.
2.1. Preliminaries of Model Watermarking
Watermarking is an IP protection technique for identifying piracy models. Specifically, a piracy model is an unauthorized copy or finetuned version of a protected AI model, exploited by adversaries for their purposes. The model watermarking consists of two stages, i.e., watermark injection and watermark verification. In the watermark injection phase, unique secret keys $\mathcal{K}$ and corresponding IP message $\mathcal{I}$ will be built and injected into the deep model $f$. As Fig. 1 shows, a typical form of secret keys and IP messages for a supervised GNN classifier are special input samples and corresponding model predictions. When verifying the ownership of model ${f}$, the model owner will extract the IP message with $\mathcal{K}$ by ${\mathcal{I}}=ExtractIP({f},\mathcal{K})$. A piracy model will retrain an IP message similar to the protected model, while an independent model will not.
2.2. Threat Model
In this paper, we focus on the IP protection of pretrained GNN encoders. Adversaries will illegitimately deploy the protected GNN encoder $f_{E}$ to obtain the classifier ${f}_{A}:\mathcal{G}\rightarrow y$ for their target classification tasks. Parameters of protected GNN encoders may be finetuned by adversaries. The target tasks of adversaries are unknown to the model owner during the watermark injection phase. During the verification phase, we assume a black-box setting, where the model owner can only obtain the predictions of the suspect model on queried samples. This is reasonable because adversaries are typically reluctant to disclose models constructed without authorization.
2.3. Problem Definition
In this paper, we focus on pretraining for graph-level tasks. Let $\mathcal{D}=\{\mathcal{G}_{i}\}_{i=1}^{|\mathcal{D}|}$ denote the set of unlabeled graphs used for GNN encoder pretraining. Given the above description, the problem of watermarking the pretraining of GNNs can be formulated by:
@@ 2.3. Problem Definition
Given a training set $\mathcal{D}$, we aim to construct and inject $(\mathcal{K},\mathcal{I})$, i.e., secret keys and IP message, into the pretraining of GNN encoder $f_{E}:\mathcal{G}\rightarrow\mathbb{R}^{d}$. The watermarked GNN encoder $f_{E}$ should produce high-quality embeddings effectively in downstream tasks. And when adversaries deploy and finetune watermarked GNN encoder $f_{E}$ to build unauthorized classifier ${f}_{A}:\mathcal{G}\rightarrow y$, the IP message can be preserved and extracted from the piracy model ${f}_{A}$ by $\mathcal{I}=ExtractIP({f}_{A},\mathcal{K})$. ${f}_{A}$ is composed of a classifier $f_{C}$ on top of $f_{E}$.
Figure 2. Overview framework of our PreGIP.
3. Methodology
@@
In this section, we give the details of our proposed PreGIP, which is illustrated in Fig. 2. As shown in Fig. 2, PreGIP aims to enforce similar representations for multiple pairs of watermark graphs. Then, downstream classifiers topped on the pretrained GNN encoder will give the same predictions on each pre-defined watermark graph pair. As a result, the consistency of predictions on the paired watermark graphs can indicate whether the suspect classifier is constructed from the watermarked GNN encoder. There are two technical challenges that remain to be addressed: (i) how can we watermark the representation space of the pretrained GNN encoder and simultaneously preserve the high quality of the learned representation space? (ii) adversaries may finetune the pre-trained GNN for various downstream tasks. How to ensure the watermark is preserved after finetuning the pretrained GNN encoder by adversaries? To address the above challenges, PreGIP incorporates a task-free watermarking loss that guarantees both watermarking performance and discriminability of graph embeddings. A finetuning-resistant watermark injection algorithm is also deployed in PreGIP. Next, we give the details of each component.
3.1. Task-Free Watermarking Framework
In this subsection, we give the details of the task-free watermarking framework for IP protection on pretraining GNNs.
3.1.1. Overview of Framework
@@
Since adversaries’ downstream tasks are generally unavailable for the pretrained GNN owner, we propose a task-free watermarking framework. As presented in Fig. 2, multiple pairs of watermark graphs $\{(\mathcal{G}_{w}^{a},\mathcal{G}_{w}^{b})\}_{w=1}^{|\mathcal{K}|}$ are constructed as secret keys $\mathcal{K}$ for IP protection. During the watermark injection phase, the proposed task-free watermark loss will ensure highly similar representations for each pair of watermark graphs $(\mathcal{G}_{w}^{a},\mathcal{G}_{w}^{b})$, while preserving the quality of representation space. In Theorem 3.1, we theoretically show that similar representations will yield the same predictions from the downstream classifier. Thus, the consistency of predictions from the suspect GNN classifier $f_{A}$ on each pair of $(\mathcal{G}_{w}^{a},\mathcal{G}_{w}^{b})$ can work as an IP message to identify unauthorized deployment of watermarked GNN encoder. Next, we present the detailed implementation of task-free watermarking.
3.1.2. Watermark Graph Pair Construction
@@
To watermark pretraining of the GNN encoder, PreGIP will enforce similar representations on distinct graphs, leading to identical predictions in the case of the piracy model. One straightforward way of construction is to randomly sample real graphs as watermark graph pairs. However, enforcing close embeddings for distinct real graphs may significantly degrade the embedding quality on real-world test samples. Therefore, we utilize synthetic graphs to construct the watermark graph pair set $\mathcal{K}=\{(\mathcal{G}_{w}^{a},\mathcal{G}_{w}^{b})\}_{w=1}^{|\mathcal{K}|}$. Specifically, the node attributes $\mathbf{x}$ of each graph in $\mathcal{K}$ are sampled by:
(1) $\mathbf{x}\sim\mathcal{N}(\bf{\mu},\bf{\sigma}^{2}),$
@@
where $\mu$ and $\sigma$ denote the mean and standard deviation of the node attributes in the pretraining dataset $\mathcal{D}$, respectively. As for the graph structure generation, we use an Erdos-Renyi (ER) random graph model (ERDdS and R&wi, 1959) which independently connects each pair of nodes with probability $p$. Note that for each pair of watermark graphs $(\mathcal{G}_{w}^{a},\mathcal{G}_{w}^{b})$, their node features and graph structures are independently synthesized, resulting in two distinct graphs. Additionally, the probability parameter $p$ and the size of the graph will be set as different values to further make $\mathcal{G}_{w}^{a}$ and $\mathcal{G}_{w}^{b}$ distinct. In this way, $\mathcal{G}_{w}^{a}$ and $\mathcal{G}_{w}^{b}$ are ensured to receive dissimilar predictions from a model independently trained on the non-watermarked dataset.
3.1.3. Watermarking Loss
@@
Apart from the term for high similarity between paired watermark graphs, we also incorporate a loss term to encourage that embeddings of synthetic watermark graphs have no overlap with those of real-world graphs. This will offer two advantages over the loss that solely enforces similar embeddings between watermark graph pairs: (i) it prevents the trivial solution that $f_{E}$ outputs the same embeddings to all graphs; (ii) embedding space of synthetic watermark graphs will remain distinct from that of real-world graphs, thereby preserving the quality of real graph embeddings for downstream tasks.
@@
Specifically, let $\mathbf{h}_{w}^{a}$ and $\mathbf{h}_{w}^{b}$ denote the graph representations of each pair of watermark graphs $(\mathcal{G}_{w}^{a},\mathcal{G}_{w}^{b})$ generated by $f_{E}$, the watermarking loss is formulated as:
(2) $\min_{\theta}$ $\quad\mathcal{L}_{W}(\theta)=\sum_{(\mathcal{G}_{w}^{a},\mathcal{G}_{w}^{b})\in\mathcal{K}}\Big{(}\|\mathbf{h}_{w}^{a}-\mathbf{h}_{w}^{b}\|^{2}+$
@@
(3) $\min_{\theta}\quad\mathcal{L}_{Pre}(\theta)+\lambda\mathcal{L}_{W}(\theta),$
@@
3.1.4. Ownership Verification
For a piracy classifier that utilizes the watermarked GNN encoder $f_{E}$, it will be more likely to give identical predictions on each pair of $(\mathcal{G}_{w}^{a},\mathcal{G}_{w}^{b})$; while an independently trained classifier would not exhibit this behavior. Thus, we extract the IP message by computing the portion of the same predictions on the watermark graph pair set $\{(\mathcal{G}_{w}^{a},\mathcal{G}_{w}^{b})\}_{w=1}^{|\mathcal{K}|}$. Given a GNN classifier $f_{A}:\mathcal{G}\rightarrow y$ to be verified, its IP indication score for ownership verification is:
(4) $I(f_{A})=\frac{\sum_{w=1}^{|\mathcal{K}|}\mathbbm{1}(f_{A}(\mathcal{G}_{w}^{a})=f_{A}(\mathcal{G}_{w}^{b}))}{|\mathcal{K}|},$